y varies directly as x^2 . y is 24 when x is 4. Find y when x is 2
Answer by stanbon(48564)

y varies directly as x^2
y = k*x^2
---------------
y is 24 when x is 4
Use this information to find "k":
24 = k/16
k = 384
--------------
EQUATION:
y = 384x^2
Find y when x is 2
y = 384*4
y = 1536
-------------
Cheers,
Stan H.